the sum of 2 numbers is 39 and the product is 360 what is the formula to calculate. I used my calculator and got 24 and 15 but my math teacher says I shouldn't use trial and error method

Answer by ankor@dixie-net.com(22740)

the two numbers: x, y
:
Write an equation for each statement:
:
the sum of 2 numbers is 39
x + y = 39
y = (39-x); use this form for substitution
:
the product is 360
x * y = 360
Substitute (39-x) for y
x * (39-x) = 360
39x - x^2 = 360
A quadratic equation
-x^2 + 39x - 360 = 0
;
what is the formula to calculate.
Solve the above equation for x, then find y